Compute the sulfide-ion concentration of a 0.5 F HC2H3O2 solution which has been saturated with H2S at 30°C.


1
Expert's answer
2021-03-29T05:15:21-0400

Constant Ionization acetic acid = 1.8×10^-5

0.5 aqueous solution of H3O+ ions concentration = 0.5 ×10^-3 mol/L

Ionization constant of H2S

1.26 × 10^-7 = 1.34 × 10^-3 × (x/0.5)

6.3 × 10^ -8 = (1.34 × 10^ -3)x

x= 6.3× 10^-8 /1.34 ×10^-3

= 4.7×10^-5 mol/L

1.26 × 10^-7 = 1.34 × 10^-3 + 4.7 × 10^-5) × x/0.5

x = 4.6 × 10 ^ -6 mol/L